OpenStudy (anonymous):

PLEASE HELP WILL GIVE MEDAL!!!!!! Which of the ordered pairs in the form (x, y) is a solution of this equation? 5x-y/3=13 (2,-9),(3,-6) A. The first is not a solution, but the second is. B. Neither is a solution. C. The first is a solution, but the second is not. D. Both are solutions

OpenStudy (er.mohd.amir):

put x=2 and y=-9 then check again put x=3 y=-6 check first 5x-y/3=13 5*2-(-9)/3=10+3=13 similarly check second.(2,-9) is solution